OpenStudy (anonymous):

verify the identity. cot (x-pi/2) = -tan x

OpenStudy (anonymous):

I'm so stumped on this one. I can't figure it out at all, and I couldn't find any problems like it that might've been asked in the past.... lol

rvc (rvc):

do you know the formula of cot(A-B) ?

rvc (rvc):

\(\huge cot(A-B)=\bf \frac{cotAcotB+1}{cotB-cotA}\)

rvc (rvc):

@zeeke u here?

OpenStudy (anonymous):

Ah sorry, I was out for a bit. I assumed that's how it would begin, but I got a bit confused by what happens afterwards.

OpenStudy (anonymous):

(cot x cot pi/2 + 1)/ (cot pi/2 - cot x)?

OpenStudy (welshfella):

cot pi/2 = 0 so if we plug this value into the above formula we get (0 +1) / (0 - cotx) = -1 cot x = -tan x

OpenStudy (anonymous):

@welshfella oh perfect! Thank you! I gave you a medal! :)

OpenStudy (welshfella):

ty yw